# Imprints of Supermassive Black Hole Evolution on the Spectral and Spatial Anisotropy of Nano-Hertz Stochastic Gravitational-Wave Background

### v1

@@ 6.2 Simulation Methodology
We have constructed all the parts required to calculate the larger-scale observation of the GW background signal from the sub-pc scale SMBHBs. The sub-pc scale physics provides us with the number distribution of binary separations ($a$) of the sources, which determines the occupation fraction of galaxies with those separations. The ROMULUS simulation informs us about the masses of the SMBHBs that reside in galaxies with a given stellar mass. The catalog provides the number density of galaxies with a given stellar mass. The relationship between the SMBH mass and the galaxy stellar mass, combined with the galaxy catalog, allows us to determine the population of SMBHBs and, consequently, its large-scale distribution. The anisotropic SGWB can be determined by calculating the GW signal from individual sources. Combining all these terms, we can express the term $\frac{d^{n+2}N(\hat{n})}{d\Theta_{n}dVdf_{r}}$ in Eq. 4 as
$\frac{d^{n+2}N(\hat{n})}{d\Theta_{n}dVdf_{r}}=$ $\frac{d^{5}N(\hat{n})}{d\log(M_{*})d\log(M_{\rm BH})dqdVdf_{r}}$ (18)
$\propto$ $\frac{dN(\hat{n})}{d\log(M_{*})}\times P_{o}(M_{*})$
×$\times P(M_{\rm BH}|M_{*},z)\times P(q|M_{*},z)\times P(f_{r}|M_{*},z)$
@@
Figure 7: SMBH mass-Stellar mass : (a) $\Omega_{\rm gw}$(f) for different $\eta$ and $\rho$ (b) $C_{l}$ vs Frequency for different $\eta$ and $\rho$, for $\ell$ = 15.
@@
Figure 8: SMBH mass-Stellar mass : (a) $\Omega_{\rm gw}$(f) for different $\nu$ (b) $C_{l}$ vs Frequency for different $\nu$, for $\ell$ = 15.
@@
Figure 9: Environmental effect : (a) $\Omega_{\rm gw}$(f) for different $\kappa$ (b) $C_{l}$ vs Frequency for different $\kappa$, for $\ell$ = 15.
@@
Figure 10: Coalescence rate evolution : (a) $\Omega_{\rm gw}$(f) for different $\delta_{f}$, $\mu$ and $z_{0}$ (b) $C_{l}$ vs Frequency for different $\mu$ and $z_{0}$, for $\ell$ = 15.
@@
We calculate the SGWB by Monte Carlo sampling of the population using eq. 18 and adding the contribution from each sampled source using eq. 4. We first assign the component masses of the binary to the host galaxies. For this, we use MICECAT galaxy catalog. We simulate the background for different SMBH mass-galaxy stellar mass (SMBH-Stellar mass) relations and different host galaxy environmental scenarios. Simulations by (Saeedzadeh et al., 2023a) have shown that most of the SMBHBs important for the PTA band reside in galaxies with high stellar mass and low star formation rate (SFR). The SMBHBs with mass $>10^{8}M_{\odot}$ are more likely to be found in the galaxies with stellar mass above $10^{10.5}M_{\odot}$ Saeedzadeh et al. (2023a). We select the host galaxies of MICECAT with stellar mass ($\rm M_{*}$) greater than that corresponding to the $\rm M_{\mu}=10^{8}M_{\odot}$ in eq. 10.
@@
We then sample a primary BH mass ($\rm M_{\rm BH}$) using eq. 10 and the mass ratio q using eq. 11 and obtain different Monte Carlo realizations of the SGWB corresponding to different realizations of SMBHBs distribution. If the SGWB signal is driven by a large number of sources (as in the lower frequencies, see Fig. 6), then the distribution is Gaussian, and the use of mode, median, and mean represents the underlying distribution. However, for higher frequencies, we have fewer sources (see Fig. 6), resulting in a non-Gaussian distribution of the SGWB signal. In this case, the mode is a better representative of the underlying distribution than the mean or median. The Monte Carlo realizations of the SGWB are normalized such that the mode of the distribution agrees with the data of the n-Hz signal observed by PTA. As a mode of distribution captures the most likely value of the underlying population (even for a non-Gaussian distribution), we expect the observed realization of SGWB of the Universe to be a typical distribution for a given underlying population model of the GW source. So, we have chosen the mode of the distribution as opposed to the mean or median in this analysis.
@@
We select sources from the sample with equal probability, ensuring that the mode of the distribution of $\Omega_{gw}$(f) in each frequency bin agrees with the current estimate from the NANOGrav 15-year data release (Agazie et al., 2023a, b). The details of the fitting can be found in the Appendix. A. In Table. 1, we show different cases for the SMBH population models considered in this work along with the parameters of the fitted power law curve (B×$\times(f/2\times 10^{-9})^{\lambda}$ ) to the mode of the $C_{l}(f)$. In the first two cases (SMBH mass-Stellar mass and SMBH mass-Stellar mass evolution), we normalize the mode of the $\Omega_{\rm gw}(f)$ by the 1/8 th of 50 % confidence level of the power law curve parameters A and $\alpha$ fitted to the 15-year data release of NANOGrav. The factor 1/8 is taken to match the 1/8 part of the sky spanned by the MICECAT catalog. In the third and fourth cases (Environmental effect and Coalescence rate), the quantity of interest is the frequency distribution, i.e., the shape of the $\Omega_{gw}(f)$. Therefore, we fix the total number of binaries to be equal to that in the third scenario of the SMBH mass-Stellar mass case (which is obtained after normalizing to the PTA data), instead of normalizing the $\Omega_{\rm gw}(f)$ by NANOGrav data again.
@@
8.5 1 0 0 5 ×$\times 10^{-9}$ 0.4 1.36 0.48
$M_{\rm BH}$ and $M_{*}$ relation 8 1.2 0 0 5 ×$\times 10^{-9}$ 0.4 0.98 0.49
8 1 0 0 5 ×$\times 10^{-9}$ 0.4 0.75 0.53
8 1 0.5 0 5 ×$\times 10^{-9}$ 0.4 1.1 0.62
Redshift evolution of $M_{\rm BH}$ and $M_{*}$ relation 8 1 0 0 5 ×$\times 10^{-9}$ 0.4 0.75 0.53
8 1 -0.5 0 5 ×$\times 10^{-9}$ 0.4 0.92 0.43
Environmental effect 8 1 0 0 1 0 0.8 5 ×$\times 10^{-9}$ 0.4 0.65 0.71
8 1 0 0 1 3 0.8 5 ×$\times 10^{-9}$ 0.4 0.66 0.63
8 1 0 1 1 1 3 0.8 5 ×$\times 10^{-9}$ 0.4 0.42 1.13
Coalescence rate 8 1 0 1 0.5 1 3 0.8 5 ×$\times 10^{-9}$ 0.4 0.47 0.94
8 1 0 0 1 3 0.8 5 ×$\times 10^{-9}$ 0.4 0.66 0.63
8 1 0 0 0 3 0.8 5 ×$\times 10^{-9}$ 0.4 0.52 0.96
Table 1: Table representing all the cases of SMBHB population considered in the work. The presence of dash for parameters $\delta$, $\kappa$ and $\gamma$ in certain cases implies that the mode of the $\Omega_{gw}$ is normalized by 50 % confidence level of the power law curve parameters (A and $\alpha$) fitted to the 15-year data release of NANOGrav. The value of the fitted parameter B and $\lambda$ of a power law to the mode of the $C_{\ell}(f)$ (over 1000 realizations) vs F is also shown.
@@ 7 Results: Isotropic and Anisotropic SGWB from simulations
Quantifying the n-Hz SGWB signal using both monopole power spectrum $\Omega_{\rm gw}(f)$ and angular power spectrum $C_{l}(f)$ as a function of frequencies brings complementary information. The isotropic SGWB signal $\Omega_{\rm gw}(f)$ measures the all-sky-averaged signal as a function of frequency, representing the net contribution across the observable Universe. However, the amplitude of the signal cannot distinguish between the scenarios on whether the signal is arising from a few very massive SMBHBs present at low redshift, or a large number of moderately massive SMBHBs spanning from very high redshift. Also, it cannot explore the underlying correlation with the galaxy properties. On the other hand, spatial anisotropy in SGWB measured in terms of $C_{l}(f)$ captures the spatial distribution of the SMBHBs. If the SMBHs form efficiently in the high redshift Universe and merge efficiently, the number of sources will be higher, and the spatial anisotropy will be less pronounced. In contrast, if the SMBHs formation is not efficient at high redshift, and they also evolve and merge inefficiently, then the spatial anisotropy will be pronounced. Moreover, as the number distribution of galaxies in heavier halos is more at low redshift than at the higher redshifts, if the host galaxies of the SMBHs are arising from heavier halos than the lighter halos, then the corresponding redshift distribution of the SMBHs will be more at low redshift. This will also exhibit a different spatial distribution of the SMBHBs and hence a different spatial distribution of the n-Hz signal. We describe below a few cases explored in this work to show the impact of SMBHB evolution on the spatial-spectral anisotropy of the SGWB signal.
@@
In Fig. 7, we show the mode of $\Omega_{\rm gw}(f)$ (over 1000 realizations) vs frequency for different scenarios, along with a 68 % confidence interval around the mode, for different SMBH mass-Stellar mass relations. We present the mode of the $\Omega_{gw}(f)$ as well as $C_{\ell}(f)$ instead of the mean, as the distributions of both quantities are highly skewed. MICECAT only covers one-eighth of the sky, so the magnitude of $\Omega_{gw}(f)$ is one-eighth that of actual all sky $\Omega_{gw}(f)$. The parameter $\eta$ captures the minimum mass of the galaxies needed to host the SMBHBs that are important for PTA sensitivity, while $\rho$ represents the slope of the relation. A curve characterized by high $\eta$ and $\rho$ values indicates a scenario in which galaxies host heavier SMBHs. The number of sources required to achieve a comparable SGWB density is small for these cases. As a consequence, the angular power spectrum ($C_{\ell}(f)$) for these cases is expected to be larger. This can be seen in Fig. 7 where we show the mode of the $C_{\ell}(f)$ vs frequency curves at $\ell$ = 15, along with a 68 % confidence interval around the mode, for different values of the parameters $\eta$ and $\rho$. It is to be noted that the anisotropy is expected to remain similar at the scales that can be resolved by the PTA as we will show in a later section. We therefore only show the result for single $\ell$. The 68 % error bars on the $C_{\ell}(f)$ show the variation in the theoretical signal due to variation in the realizations. This is the theoretical uncertainty associated with cosmic variance due to a single realization of our Universe (Roebber et al., 2016; Pacilio et al., 2023; Allen, 2023).
@@
To understand the behavior of the anisotropy as a function of frequency, we infer the scaling of the anisotropy with frequency by a power law $C_{l}(f)=B(l)\times$ $(f/2\times 10^{-9})^{\lambda}$ to the mode of $C_{\ell}(f)$. The value of parameters can be inferred from Table 1. The power-law index for these cases is similar, with a slightly higher value for $\eta=8$ and $\rho=1$ compared to the other two cases. It is important to point out that, although all these cases have similar $\Omega_{\rm gw}(f)$, their $C_{l}(f)$ values differ across frequencies. The correlation between the masses of SMBHBs and their host galaxies provides valuable insight into the evolution and initial seed masses of SMBHs. In cases where $\eta$ is comparatively large, similar galaxies will host relatively larger SMBHs, suggesting that SMBHs originated from large seed black holes. This leads to a relatively larger anisotropy in the SGWB as discussed in the last paragraph. Similarly, a large value of $\rho$ suggests a stronger correlation of stellar mass with SMBH mass. A large value of $\rho$, indicating a strong correlation between stellar mass and SMBH mass, will result in a higher anisotropy.
@@
Redshift evolution of SMBH mass and Stellar mass relation: In Fig. 8, we depict the mode of $\Omega_{gw}(f)$ and $C_{\ell}(f)$ (more than 1000 realizations) vs. frequency, along with a confidence interval 68 % around the mode for different scenarios of evolution of the $M_{\rm BH}-M_{*}$ relation with redshift, parameterized by $\nu$). The anisotropy ($C_{\ell}(f)$) is larger for the case where $\nu$ is positive. However, the anisotropy is comparable for the cases of $\nu=0$ and $\nu=-0.5$. A positive value of $\nu$ represents a scenario where galaxies at high redshift host heavy SMBHBs, whereas a negative value of $\nu$ represents a scenario where galaxies at high redshift host lighter SMBHBs (see Fig. 4 and Fig. 4 for comparing their distribution). If heavier binaries occupy the center of a galaxy at high redshifts, a smaller number of binaries is required to generate the same SGWB. The anisotropy for cases with a negative value of $\nu$ will naturally be lower, as a larger number of binaries are required to generate the same SGWB compared to cases with a positive value of $\nu$, where we have heavier binaries at higher redshifts. The anisotropy is not very distinct between the cases when $\nu=0$ and $\nu=-0.5$, despite there being a larger number of sources for $\nu=-0.5$. This is because lighter sources at higher redshifts in the case $\nu=-0.5$ contribute very little to the overall anisotropy. The power-law index of the curve fitted to $C_{\ell}(f)$, represented by $B(\ell)\times(f/2\times 10^{-9})^{\lambda}$, tends to be slightly higher for positive values of $\nu$, as expected.
@@
It should be noted that there is a sharp jump in the $C_{\ell}(f)$ curve at $f=2.6\times 10^{-8}$ Hz for $\nu=0.5$. This is because, at high-frequency bins, the expectation value of the number of sources in the low redshift bins is less than one (as can be seen for some cases in Fig. 6). However, Monte Carlo sampling of the population mostly results in either one source or none, not a fraction of a source. Therefore, the sporadic sampling at high-frequency bins leads to a sudden increase in the anisotropy, or $C_{\ell}(f)$, at these frequencies.
@@
Figure 11: Color matrix plot of the mode of $C_{\ell}(f)$ (more than 1000 realizations) as a function of frequency (F) and multipole moment ($\ell$) for the first scenario of (a) the SMBH mass-Stellar mass case, (b) the SMBH mass-Stellar mass evolution case, (c) the case of environmental effect and (d) the case of coalescence rate.
@@
The correlation between the masses of SMBHBs and their host galaxies and the redshift evolution of the correlation offers valuable insights into the evolution as well as the initial seed masses of SMBHs. If SMBHBs are predominantly found in galaxies with very high stellar masses, this suggests that the SMBHs have undergone multiple merger cycles. This information provides clues about the initial seed masses required for the formation of these SMBHs. A small initial seed mass would require numerous merger cycles and efficient accretion processes throughout its lifetime. The redshift evolution of the $M_{\rm BH}-M_{*}$ relationship can potentially tell us about the accretion and merger history of the SMBHs. If the SMBHs are continuously growing even at low redshifts, we expect to find heavier BHs at low redshifts that reside in similar galaxies than at higher redshifts. This relates to the case with negative $\nu$ where $C_{\ell}(f)$ or anisotropy is greater than the case where $\nu$ is positive.
@@
Environmental effect : The residence time of a binary at a high frequency, in the absence of environmental effects, is proportional to $f^{-11/3}$. This relationship implies that, compared to sources inspiraling at the lower limit of the PTA band, we can expect the number of sources observable at the higher frequency band of the PTA to be extremely low. In the presence of environmental effects, the residence time of the binary decreases at lower frequencies. This is because these effects are only significant at large separations of the binary. In a scenario where the galactic environment is more efficient at driving binaries to smaller separations, we would expect to observe a population with a relatively smaller number of binaries at low frequencies than one would expect in the absence of environmental effects.
@@
In Fig. 9, we illustrate the mode of $\Omega_{gw}(f)$ and $C_{\ell}(f)$ (over 1000 realizations) as a function of frequency for different environmental effects captured by the parameter $\kappa$. Additionally, we depict a 68 % confidence interval around the mode. In these instances, we maintain a constant total number of binaries. The curves representing $C_{\ell}(f)$ at different $\kappa$ values are very similar to each other with almost the same amplitude B and very little difference in the power law index $\lambda$ of the fitted parameters to the $C_{\ell}(f)$. This is because the environmental effect is important only at low. At these lower frequencies, the number of sources is much larger, resulting in low anisotropy for both cases. This means that the effect of the spectral change due to the environmental effect is very small on a large-scale distribution of the SGWB.
@@
Coalescence rate: At last, in Fig. 10, we illustrate the result of $\Omega_{gw}(f)$ and $C_{\ell}(f)$ for different coalescence rates and their evolution. We show the mode of the $\Omega_{gw}(f)$ and $C_{\ell}(f)$ (over 1000 realizations) along with a 68 % confidence interval around the mode. The parameter $\delta_{f}$ is the power index of the coalescence rate, and parameters $\mu$ and $z_{0}$ capture the redshift evolution of the coalescence rate. In the case where $\mu$ is non-zero, we have a redshift-dependent coalescence rate. The parameter $z_{0}$ represents the redshift below which the coalescence rate of the binaries (see Eq. 16) is higher at lower frequencies. For $z_{0}=1$, the coalescence rate is higher at lower frequencies up to a redshift of 1. As the value of $z_{0}$ increases, the number of binaries emitting at high frequencies decreases at lower redshifts.

### v2

@@ 5.2 Simulation Methodology
We have constructed all the parts required to calculate the larger-scale observation of the GW background signal from the sub-pc scale SMBHBs. The sub-pc scale physics provides us with the number distribution of binary separations ($a$) of the sources, which determines the occupation fraction of galaxies with those separations. The ROMULUS simulation informs us about the masses of the SMBHBs that reside in galaxies with a given stellar mass. The MICECAT catalog provides the number density of galaxies with a given stellar mass. The relationship between the SMBH mass and the galaxy stellar mass, combined with the galaxy catalog, allows us to determine the population of SMBHBs and, consequently, its large-scale distribution. The anisotropic SGWB can be determined by calculating the GW signal from individual sources. Combining all these terms, we can express the term $\frac{d^{n+4}N(\hat{n})}{d\Theta_{n}dVdf_{r}}$ in Eq. 4 as Eq. 17
@@
Figure 6: SMBH mass-Stellar mass : (a) $\Omega_{\rm gw}$(f) for different values of $\eta$ and $\rho$ (b) $C_{l}(f)$ vs Frequency ($f$) for different values of $\eta$ and $\rho$, for $\ell$ = 15.
@@
Figure 7: SMBH mass-Stellar mass evolution : (a) $\Omega_{\rm gw}$(f) for different values of $\nu$ (b) $C_{l}(f)$ vs Frequency ($f$) for different values of $\nu$, for $\ell$ = 15.
@@
Figure 8: Environmental effect : (a) $\Omega_{\rm gw}$(f) for different values of $\kappa$ (b) $C_{l}(f)$ vs Frequency ($f$) for different values of $\kappa$, for $\ell$ = 15.
@@
Figure 9: Coalescence rate : (a) $\Omega_{\rm gw}$(f) for different values of $\delta_{f}$, $\mu$ and $z_{0}$ (b) $C_{l}(f)$ vs Frequency ($f$) for different values of $\delta_{f}$, $\mu$ and $z_{0}$, for $\ell$ = 15.
@@
We calculate the SGWB by Monte Carlo sampling of the population using eq. 17 and adding the contribution from each sampled source using eq. 4. We first assign the component masses of the binary to the host galaxies. For this, we use MICECAT galaxy catalog. We simulate the background for different SMBH mass-galaxy stellar mass (SMBH-Stellar mass) relations and different host galaxy environmental scenarios. Simulations by (Saeedzadeh et al., 2023a) have shown that most of the SMBHBs important for the PTA band reside in galaxies with high stellar mass and low star formation rate (SFR). The SMBHBs with mass $>10^{8}M_{\odot}$ are more likely to be found in the galaxies with stellar mass above $10^{11}M_{\odot}$ (Saeedzadeh et al., 2023a). We only select the host galaxies of MICECAT with stellar mass ($\rm M_{*}$) greater than that corresponding to the $\rm M_{\mu}=10^{8}M_{\odot}$ in eq. 12.
@@
We then sample a primary BH mass ($\rm M_{\rm BH}$) using eq. 10 and the mass ratio q using eq. 11 and obtain different Monte Carlo realizations of the SGWB corresponding to different realizations of SMBHBs distribution. If the SGWB signal is driven by a large number of sources (as in the lower frequencies, see Fig. 5), then the distribution is Gaussian, and the use of mode, median, and mean represents the underlying distribution. However, for higher frequencies, we have fewer sources (see Fig. 5), resulting in a non-Gaussian distribution of the SGWB signal. In this case, the mode is a better representative of the underlying distribution than the mean or median. In order to obtain the total number of SMBHBs occupying the galaxies, we normalize the realizations of the SGWB such that the mode of the distribution agrees with the data of the n-Hz signal observed by PTA. As a mode of distribution captures the most likely value of the underlying population (even for a non-Gaussian distribution), we expect the observed realization of SGWB of the Universe to be a typical distribution for a given underlying population model of the GW source. So, we have chosen the mode of the distribution as opposed to the mean or median in this analysis.
@@
We select sources from the sampled sources with equal probability, ensuring that the mode of the distribution of $\Omega_{\rm gw}$(f) in each frequency bin agrees with the current estimate from the NANOGrav 15-year data release (Agazie et al., 2023a, b). The details of the fitting can be found in the Appendix. A. In Table. 1, we show different cases for the SMBH population models considered in this work along with the parameters of the fitted power law curve (B×$\times(f/2\times 10^{-9})^{\lambda}$ ) to the mode of the $C_{\ell}(f)$. In the first two cases (SMBH mass-Stellar mass and SMBH mass-Stellar mass evolution), we choose the number of sources contributing to the signal in each frequency bin such that it is comparable to the 1/8 th power law curve fitted to the 15-year data release of NANOGrav. The factor 1/8 is taken to match the 1/8 part of the sky spanned by the MICECAT catalog. In the third and fourth cases (Environmental effect and Coalescence rate), the quantity of interest is the frequency distribution, i.e., the shape of the $\Omega_{\rm gw}(f)$. Therefore, we fix the total number of binaries to be equal to that in the third scenario of the SMBH mass-Stellar mass case (which is obtained after normalizing to the PTA data), instead of normalizing the $\Omega_{\rm gw}(f)$ by NANOGrav data again.
@@
8.5 1 0 0 5 ×$\times 10^{-9}$ 0.4 0.17 0.48
$M_{\rm BH}$ and $M_{*}$ relation 8 1.2 0 0 5 ×$\times 10^{-9}$ 0.4 0.122 0.49
8 1 0 0 5 ×$\times 10^{-9}$ 0.4 0.094 0.53
8 1 0.5 0 5 ×$\times 10^{-9}$ 0.4 0.1375 0.62
Redshift evolution of $M_{\rm BH}$ and $M_{*}$ relation 8 1 0 0 5 ×$\times 10^{-9}$ 0.4 0.094 0.53
8 1 -0.5 0 5 ×$\times 10^{-9}$ 0.4 0.115 0.43
Environmental effect 8 1 0 0 1 0 0.8 5 ×$\times 10^{-9}$ 0.4 0.0813 0.71
8 1 0 0 1 3 0.8 5 ×$\times 10^{-9}$ 0.4 0.66 0.0825
8 1 0 1 1 1 3 0.8 5 ×$\times 10^{-9}$ 0.4 0.0525 1.13
Coalescence rate 8 1 0 1 0.5 1 3 0.8 5 ×$\times 10^{-9}$ 0.4 0.059 0.94
8 1 0 0 1 3 0.8 5 ×$\times 10^{-9}$ 0.4 0.0825 0.63
8 1 0 0 0 3 0.8 5 ×$\times 10^{-9}$ 0.4 0.065 0.96
Table 1: Table representing all the cases of SMBHB population considered in the work. The presence of dash for parameters $\delta$, $\kappa$ and $\gamma$ in certain cases implies that the mode of the $\Omega_{\rm gw}$ is normalized by 50 % confidence level of the power law curve parameters (A and $\alpha$) fitted to the 15-year data release of NANOGrav. The value of the fitted parameter B and $\lambda$ of a power law to the mode of the $C_{\ell}(f)$ (over 1000 realizations) vs $f$ is also shown.
@@ 6 Results: Isotropic and Anisotropic SGWB from simulations
Quantifying the n-Hz SGWB signal using both monopole power spectrum $\Omega_{\rm gw}(f)$ and angular power spectrum $C_{\ell}(f)$ as a function of frequencies brings complementary information. The isotropic SGWB signal $\Omega_{\rm gw}(f)$ measures the all-sky-averaged signal as a function of frequency, representing the net contribution across the observable Universe. However, the amplitude of the signal cannot distinguish between the scenarios on whether the signal is arising from a few very massive SMBHBs present at low redshift, or a large number of moderately massive SMBHBs spanning from very high redshift. Also, it cannot explore the underlying correlation with the galaxy properties. On the other hand, spatial anisotropy in SGWB measured in terms of $C_{\ell}(f)$ captures the spatial distribution of the SMBHBs. If the SMBHs form efficiently in the high redshift Universe and merge efficiently, the number of sources will be higher, and the spatial anisotropy will be less pronounced. In contrast, if the SMBH formation is not efficient at high redshifts, and they also evolve and merge inefficiently, then the spatial anisotropy is expected to be more significant. Moreover, as the number distribution of galaxies in heavier halos is higher at low redshift than at the higher redshifts if the host galaxies of the SMBHs are arising from heavier halos than the lighter halos, then the corresponding redshift distribution of the SMBHs will be more at low redshift. This will also exhibit a different spatial distribution of the SMBHBs and hence a different spatial distribution of the n-Hz signal. We describe below a few cases explored in this work to show the impact of SMBHB evolution on the spatial-spectral anisotropy of the SGWB signal. Moreover, since the number distribution of galaxies in heavier halos is higher at low redshifts compared to higher redshifts, if the host galaxies of the SMBHBs occupy heavier halos rather than lighter ones, then SMBHBs will predominantly occur in lower redshifts.
@@
In Fig. 6, we show the mode of $\Omega_{\rm gw}(f)$ (over 1000 realizations) vs frequency for different scenarios, along with a 68 % confidence interval around the mode, for different SMBH mass-Stellar mass relations. We present the mode of the $\Omega_{\rm gw}(f)$ as well as $C_{\ell}(f)$ instead of the mean, as the distributions of both quantities are highly skewed. MICECAT only covers one-eighth of the sky, so the magnitude of $\Omega_{\rm gw}(f)$ is one-eighth that of actual all sky $\Omega_{\rm gw}(f)$. The parameter $\eta$ captures the minimum mass of the galaxies needed to host the SMBHBs that are important for PTA sensitivity, while $\rho$ represents the slope of the relation. A curve characterized by high $\eta$ and $\rho$ values indicates a scenario in which galaxies host heavier SMBHs. The number of sources required to achieve a comparable SGWB density is small for these cases. As a consequence, the angular power spectrum ($C_{\ell}(f)$) for these cases is expected to be larger. This can be seen in Fig. 6 where we show the mode of the $C_{\ell}(f)$ versus frequency curves at $\ell$ = 15, along with a 68 % confidence interval around the mode, for different values of parameters $\eta$ and $\rho$. It is important to note that the anisotropic signal is expected to be dominated by shot noise, and therefore, it is expected to remain similar at the scales that can be resolved by the PTA, as we will demonstrate in a later section. We therefore only show the result for single $\ell$. The 68 % error bars on the $C_{\ell}(f)$ represent the uncertainty in the theoretical signal around the mode. This uncertainty arises because of cosmic variance. The cosmic variance results from the variation of the properties of individual sources in different realizations sampled from a given population of the SMBHB. In reality, we only have one realization of the universe. Consequently, the distribution of the properties of the source, such as the masses and number density of the source, introduces inherent uncertainty in the anisotropic signal $C_{\ell}$.
@@
To understand the behavior of the anisotropy as a function of frequency, we infer the scaling of the anisotropy with frequency by fitting a power law $C_{\ell}(f)=B(\ell)\times$ $(f/2\times 10^{-9})^{\lambda}$ to the mode of $C_{\ell}(f)$. The value of parameters can be inferred from Table 1. The power-law index for these cases is similar, with a slightly higher amplitude value for $\eta=8.5$ and $\rho=1$ compared to the other two cases. It is important to point out that, although all these cases have similar $\Omega_{\rm gw}(f)$, their $C_{\ell}(f)$ values differ across frequencies. The correlation between the masses of SMBHBs and their host galaxies provides valuable insight into the evolution and initial seed masses of SMBHs. In cases where $\eta$ is comparatively large, similar galaxies will host relatively larger SMBHs, suggesting that the SMBHs originated from large seed BHs or more efficient merger and accretion. This leads to a relatively larger anisotropy in the SGWB as discussed in the last paragraph. Similarly, a large value of $\rho$ suggests a stronger correlation of stellar mass with SMBH mass. A large value of $\rho$ will result in a higher anisotropy.
@@
Redshift evolution of SMBH mass and Stellar mass relation: In Fig. 7, we depict the mode of $\Omega_{\rm gw}(f)$ and $C_{\ell}(f)$ (over 1000 realizations) vs. frequency, along with a confidence interval 68 % around the mode for different scenarios of evolution of the $M_{\rm BH}-M_{*}$ relation with redshift, parameterized by $\nu$). The anisotropy ($C_{\ell}(f)$) is larger for the case where $\nu$ is positive. However, the anisotropy is comparable for the cases of $\nu=0$ and $\nu=-0.5$. A positive value of $\nu$ represents a scenario where galaxies at higher redshifts host heavier SMBHBs, whereas a negative value of $\nu$ represents a scenario where galaxies at higher redshifts host lighter SMBHBs (see Fig. 3 and Fig. 3 for comparing their distribution). If heavier binaries occupy the center of a galaxy at high redshifts, a smaller number of binaries is required to generate the same SGWB. The anisotropy for cases with a negative value of $\nu$ will naturally be lower, as a larger number of binaries are required to generate the same SGWB compared to cases with a positive value of $\nu$, where we have heavier binaries at higher redshifts. The anisotropy is not very distinct between the cases when $\nu=0$ and $\nu=-0.5$, despite there being a larger number of sources for $\nu=-0.5$. This is because lighter sources at higher redshifts in the case $\nu=-0.5$ contribute very little to the overall anisotropy. The power-law index ($\lambda$) and the amplitude (B) of the curve fitted to $C_{\ell}(f)$, tends to be slightly higher for positive values of $\nu$, as expected.
@@
It should be noted that there is a sharp jump in the $C_{\ell}(f)$ curve at $f=2.6\times 10^{-8}$ Hz for $\nu=0.5$. This is because, at high-frequency bins, the expectation value of the number of sources in the low redshift bins is less than one (as can be seen for some cases in Fig. 5). However, Monte Carlo sampling of the population mostly results in either one source or none, not a fraction of a source. Therefore, the sporadic sampling at high-frequency bins leads to a sudden increase in $C_{\ell}(f)$ at these frequencies.
@@
Figure 10: Color matrix plot of the mode of $C_{\ell}(f)$ (over 1000 realizations) as a function of frequency ($f$) and multipole moment ($\ell$) for the first scenario of (a) the SMBH mass-Stellar mass case, (b) the SMBH mass-Stellar mass evolution case, (c) the environmental effect case and (d) the coalescence rate case.
@@
The correlation between the masses of SMBHBs and their host galaxies as well as the redshift evolution of the correlation offers valuable insights into the evolution of SMBH population. The redshift evolution of the $M_{\rm BH}-M_{*}$ relationship can potentially tell us about the accretion and merger history of the SMBHs. If the SMBHs are continuously growing even at low redshifts, we expect to find heavier BHs at low redshifts that reside in similar galaxies than at higher redshifts. This relates to the case with negative $\nu$ where the anisotropy is smaller than the case where $\nu$ is positive.
@@
Environmental effect : The residence time of a binary at a high frequency, in the absence of environmental effects, is proportional to $f^{-11/3}$. This implies that, compared to sources inspiraling at the lower limit of the PTA band, we can expect the number of sources observable at the higher frequency band of the PTA to be extremely low. However, in the presence of environmental effects, the residence time of the binary decreases at lower frequencies. This is because these effects are only significant at large separations of the binary. In a scenario where the galactic environment is more efficient at driving binaries to smaller separations, we would expect to observe a population with a relatively smaller number of binaries at low frequencies than one would expect in the absence of environmental effects.
@@
In Fig. 8, we illustrate the mode of $\Omega_{\rm gw}(f)$ and $C_{\ell}(f)$ (over 1000 realizations) as a function of frequency for different environmental effects captured by the parameter $\kappa$. Additionally, we depict a 68 % confidence interval around the mode. In these instances, we maintain a constant total number of binaries. The curves representing $C_{\ell}(f)$ at different $\kappa$ values are very similar to each other with almost the same amplitude B and very little difference in the power law index $\lambda$ of the fitted parameters to the $C_{\ell}(f)$. This is because the environmental effect is important only at low frequencies. At these lower frequencies, the number of sources is much larger, resulting in low anisotropy for both cases. This means that the effect of the spectral change due to the environmental effect is very small on a large-scale distribution of the SGWB.
@@
Coalescence rate: At last, in Fig. 9, we illustrate the result of $\Omega_{\rm gw}(f)$ and $C_{\ell}(f)$ for different coalescence rates and their evolution. We show the mode of the $\Omega_{\rm gw}(f)$ and $C_{\ell}(f)$ (over 1000 realizations) along with a 68 % confidence interval around the mode. The parameter $\delta_{f}$ is the power index of the coalescence rate, and parameters $\mu$ and $z_{0}$ capture the redshift evolution of the coalescence rate. In the case where $\mu$ is non-zero, we have a redshift-dependent coalescence rate. The parameter $z_{0}$ represents the redshift below which the coalescence rate of the binaries (see Eq. 16) is higher at lower frequencies. For $z_{0}=1$, the coalescence rate is higher at lower frequencies up to a redshift of 1. As the value of $z_{0}$ increases, the number of binaries emitting at high frequencies decreases at lower redshifts.